in the following system of equations, what does y equal?
8x + 5y=5
4x + y=9

Answer by ewatrrr(24785)

 
Hi,
8x + 5y =5
4x + y = 9 |Multiply 2nd Eq by 2 and subtacting from the 1st to eliminate x
3y = -13
y = -13/3